A pound of tin is worth $1 more than a pound of copper. Two pounds of tin are mixed with 8 pounds of copper to make bronze that sells for $4.45 per pound. How much is a pound of tin worth?
Answer by stanbon(75887)

A pound of tin is worth $1 more than a pound of copper. Two pounds of tin are mixed with 8 pounds of copper to make bronze that sells for $4.45 per pound. How much is a pound of tin worth?
----------------
2t+8c = 445 cents
t = c + 100 cents
------------------------------
Substitute for "t" and solve for "c":
2(c+100) + 8c = 445
-----
2c + 200 + 8c = 445
10c = 245
c = 24.5 cents (price of lb of copper)
----
Solve for "t":
t = c + 100
t = 24.5 + 100
t = 124.5 cents (price of lb of tin)
--------------------------------------
Cheers,
Stan H.
=======================